# VideoAnydoor: High-fidelity Video Object Insertion with Precise Motion Control

### v1

@@ Abstract
Despite significant advancements in video generation, inserting a given object into videos remains a challenging task. The difficulty lies in preserving the appearance details of the reference object and accurately modeling coherent motion at the same time. In this paper, we propose VideoAnydoor, a zero-shot video object insertion framework with high-fidelity detail preservation and precise motion control. Starting from a text-to-video model, we utilize an ID extractor to inject the global identity and leverage a box sequence to control the overall motion. To preserve the detailed appearance and meanwhile support fine-grained motion control, we design a pixel warper. It takes the reference image with arbitrary key-points and the corresponding key-point trajectories as inputs. It warps the pixel details according to the trajectories and fuses the warped features with the diffusion U-Net, thus improving detail preservation and supporting users in manipulating the motion trajectories. In addition, we propose a training strategy involving both videos and static images with a reweight reconstruction loss to enhance insertion quality. VideoAnydoor demonstrates significant superiority over existing methods and naturally supports various downstream applications (e.g., video face swapping, video virtual try-on, multi-region editing) without task-specific fine-tuning.
@@ 1 Introduction
Faced with this challenge, we attempt to accurately preserve the object’s identity and precisely control the object’s motion throughout the whole video. Specifically, we propose an end-to-end framework termed VideoAnydoor. Starting from a text-to-video diffusion model, the concatenation of random noise, object masks, and the masked video is utilized as input. Meanwhile, the reference image with no background is fed into the ID extractor to extract compact and discriminative ID tokens. Then these ID tokens are injected into the diffusion model together with the box sequence as coarse guidance of identity and motion to generate the desired composition. Additionally, a pixel warper module is designed for joint modeling of the fine-grained appearance and precise motion. It takes the reference image with arbitrary key-points and the corresponding key-point trajectories to warp the pixel details according to the desired motion. To further enhance motion alignment with the given condition, a re-weighted reconstruction loss is introduced. Moreover, to address the scarcity of high-quality videos, we augment high-quality image data as videos to improve the alignment of appearance details. As shown in Fig. 1, with these techniques, users can edit specific regions in the video by providing target images and drawing boxes and trajectory lines. It should be noted that our inserted object is not constrained by shape or appearance, demonstrating robustness and generality in diverse scenarios.
@@
  • We design multiple strategies to further enhance the capability of accurate insertion, including image-video mix training, training trajectory filtering, reweight reconstruction losses. Extensive experiments demonstrate their effectiveness in precise ID preservation and motion control.
@@ 2 Related Work
Figure 2: The pipelines of our VideoAnydoor. First, we input the concatenation of the original video, object masks, and masked video into the 3D U-Net. Meanwhile, the background-removed reference image is fed into the ID extractor, and the obtained features are injected into the 3D U-Net. In our pixel warper, the reference image marked with key points and the trajectories are utilized as inputs for the content and motion encoders. Then, the extracted embeddings are input into cross-attentions for further fusion. The fused results serve as the input of a ControlNet, which extracts multi-scale features for fine-grained injection of motion and identity. The framework is trained with reweight reconstruction losses. We use a blend of real videos and image-simulated videos for training to compensate for the data scarcity.
@@ 3.1 Overview of Framework
The VideoAnydoor pipeline is illustrated in Fig. 2. To reconstruct the background within the masked region, we build our method on a 2D in-painting diffusion model. Specifically, following the latent diffusion model [31], we encode both the source video and the masked video with a VAE encoder to obtain the latent space representations $z_{ori}$ and $z_{mask}$. The corresponding masks are 8 times down-sampled as the mask latent. Subsequently, DDIM inversion [5] is applied to transform the clean latent $z_{ori}$ back to the noisy latent $z_{T}$. Then we concatenate $z_{T}$, $z_{mask}$ and $z_{mask}$ as a 9-channel tensor for 3D U-Net. To utilize the priors of high-quality video generation, we integrate the motion layers [11] into the in-painting model as the 3D U-Net to ensure essential temporal consistency. For coarse-grained control, we leverage the powerful visual encoder DINOv2 [25] as the ID extractor for ID preservation and use the bounding boxes as motion guidance. Before feeding the reference image into the extractor, we remove its background with a segmentor [16] and align the object to the image center to retain compact and ID-related representations. For fine-grained control, we adopt the interaction-friendly trajectory lines as the control signals and propose a pixel warper to warp the pixel details according to the desired motion for joint modeling of appearance details and precise motion. Finally, we introduce a reweight reconstruction loss to amplify the influence of key-points and design a novel image-video mix-training strategy to address the scarcity of high-quality video data.
For convenience, the trajectory map and correspondence reference image are denoted as $c_{mot}\in\mathcal{R}^{N\times 3\times H\times W}$ and $c_{ref-key}\in\mathcal{R}^{3\times H\times W}$ respectively. $V\in\mathcal{R}^{N\times 3\times H\times W}$, $M\in\mathcal{R}^{N\times 1\times H\times W}$ and $c_{ref}\in\mathcal{R}^{3\times H\times W}$ represent the original video, masks of the edited region and the reference image respectively. $N,H,W$ is the frame number, height, and width of the original video. The content and motion encoders are denoted as $E_{c}$ and $E_{m}$ respectively.
@@ 3.2 Pixel Warper
Reweight reconstruction.
To further enhance the fine-grained modeling of identity and motion, we propose a reweighted diffusion loss that differentiates the contributions of regions inside the bounding boxes and areas around the trajectories. Specifically, we amplify the contributions within bounding boxes and regions around trajectories to enhance subject and motion learning while preserving the original diffusion loss for regions outside these boxes. Denote the masks of the regions that $c_{mot}$ covers as $\mathbf{M_{mot}}$, we perform 8 times down-sampling on $\mathbf{M_{mot}}$ and $\mathbf{M}$, which are denoted as $\mathbf{M_{mot}}^{\prime}$ and $\mathbf{M^{\prime}}$ respectively. The reconstruction loss can be formulated as:
$\begin{split}\mathcal{L}(\theta)=\mathbb{E}_{\boldsymbol{z},\epsilon,\boldsymbol{c},t}[({\lambda_{\mathcal{L}}(\mathbf{M^{\prime}}}+\mathbf{M_{mot}^{\prime}})+{(1-\mathbf{M^{\prime}})})\\ \cdot\mathopen{}\mathclose{{}\left\|\epsilon-\epsilon_{\theta}\mathopen{}\mathclose{{}\left(\boldsymbol{z}_{t},\boldsymbol{c}_{\mathrm{ref}},\boldsymbol{c}_{\mathrm{mot}},\boldsymbol{c}_{\mathrm{ref-key}},t}\right)}\right\|_{2}^{2}],\end{split}$ (2)
where $\lambda_{\mathcal{L}}$ denotes the balancing loss weight. $\epsilon_{\theta},\epsilon$ are the prediction of the 3D U-Net and the target.
@@ 4.1 Experimental Setup
For comprehensive evaluation, we construct a benchmark consisting of around 200 videos collected from Pexel, which includes ten different categories (e.g., persons, dogs). We also make qualitative analysis on the ViViD [8] and CHDTF [53] test set to evaluate the performance for virtual video try-on and video face swapping.
@@ 4.3 Quantitative Comparison
Only Real-video Data 34.4 76.4 52.0
Only Static-image Data 34.1 76.2 51.2
FrozenDINOv2 33.2 74.5 51.4
w/o $\rm PixelWarper^{\dagger}$ 35.1 77.0 53.1
w/o Pixel Warper 33.6 72.1 48.1
w/o Re-weighted Loss 35.1 77.0 53.1
Ours-full 37.7 81.1 58.9
@@ 4.4 Ablation Studies
Only Real-video Data 66.1 67.0 69.6
Only Static-image Data 72.3 72.6 75.1
@@
w/o $\rm PixelWarper^{\dagger}$ 81.3 82.2 85.0
w/o $\rm PixelWarper$ 78.3 81.7 84.0
w/o Re-weighted Loss 75.4 84.2 85.1
Ours-full 88.0 91.1 92.3
@@
We first conduct an investigation of the core components on ID preservation. From Tab. 4, we can observe that training with fixed DINOv2 induces much inferior performance. Similarly, training only on videos suffers from severe accuracy degradation across all the metrics. Moreover, our pixel warper can effectively help inject the appearance details according to the motion. The results also show that our reweight reconstruction loss can bring a performance boost to the baseline by making the model focus on the foreground regions. The best performance can be achieved by training with all the modules.
@@
Random X-Pose points 80.4 82.2 82.8
Grid points 82.6 83.7 85.2
w/o NMS 82.3 83.1 84.6
Tight box 83.2 85.4 86.1
Ours-full 88.0 91.1 92.3
@@ 4.5 More Applications
Video face swapping. Besides video try-on, we further apply our method to video face swapping, which requires more precise control of tiny movements and preservation of face identities. Specifically, we conduct evaluations on the CHDTF dataset [53] and use 16 points with the largest movement in the face landmarks as the initial key-points. As demonstrated in Fig. 8, VideoAnydoor could give satisfactory performance for this task as well.
@@ 5 Conclusion
In this paper, we present VideoAnydoor for end-to-end video object insertion with precise motion control. Specifically, it can effectively characterize the reference target with an ID extractor when trained with a combination of videos and high-quality images. Moreover, it can achieve smooth motion consistency and effective preservation of appearance details through the proposed pixel warper. Our VideoAnydoor has promising performance on diverse precise video editing applications, e.g., (1) object insertion, (2) virtual video try-on, (3) video face swapping, and (4) multi-region editing. Extensive qualitative and quantitative experimental results demonstrate its superiority over previous methods. It provides a universal solution for general region-to-region mapping tasks as well.

### v3

@@ Abstract
Despite significant advancements in video generation, inserting a given object into videos remains a challenging task. The difficulty lies in preserving the appearance details of the reference object and accurately modeling coherent motion at the same time. In this paper, we propose VideoAnydoor, a zero-shot video object insertion framework with high-fidelity detail preservation and precise motion control. Starting from a text-to-video model, we utilize an ID extractor to inject the global identity and leverage a box sequence to control the overall motion. To preserve the detailed appearance and meanwhile support fine-grained motion control, we design a pixel warper. It takes the reference image with arbitrary key-points and the corresponding key-point trajectories as inputs. It warps the pixel details according to the trajectories and fuses the warped features with the diffusion U-Net, thus improving detail preservation and supporting users in manipulating the motion trajectories. In addition, we propose a training strategy involving both videos and static images with a weighted loss to enhance insertion quality. VideoAnydoor demonstrates significant superiority over existing methods and naturally supports various downstream applications (e.g., video face swapping, video virtual try-on, multi-region editing) without task-specific fine-tuning.
@@ 1 Introduction
Faced with this challenge, we attempt to accurately preserve the object’s identity and precisely control the object’s motion throughout the whole video. Specifically, we propose an end-to-end framework termed VideoAnydoor. Starting from a text-to-video diffusion model, the concatenation of random noise, object masks, and the masked video is utilized as input. Meanwhile, the reference image with no background is fed into the ID extractor to extract compact and discriminative ID tokens. Then these ID tokens are injected into the diffusion model together with the box sequence as coarse guidance of identity and motion to generate the desired composition. Additionally, a pixel warper module is designed for joint modeling of the fine-grained appearance and precise motion. It takes the reference image with arbitrary key-points and the corresponding key-point trajectories to warp the pixel details into the target regions according to the desired motion and pose. Moreover, to address the scarcity of high-quality videos, we manually augment extensive high-quality image data as videos to improve the fine-grained alignment of appearance details. As shown in Fig. 1, with these techniques, users can edit specific regions in the video by providing target images, drawing boxes and trajectory lines. It should be noted that our inserted object is not constrained by shape, appearance or the range of the given movement, demonstrating great robustness and generality in diverse scenarios.
@@
  • We design multiple strategies to further enhance the capability of accurate insertion, including image-video mix training, training trajectory filtering. Extensive experiments demonstrate their effectiveness in precise ID preservation and motion control.
@@ 2 Related Work
Figure 2: The pipelines of our VideoAnydoor. First, we input the concatenation of the original video, object masks, and masked video into the 3D U-Net. Meanwhile, the background-removed reference image is fed into the ID extractor, and the obtained features are injected into the 3D U-Net. In our pixel warper, the reference image marked with key points and the trajectories are utilized as inputs for the content and motion encoders. Then, the extracted embeddings are input into cross-attentions for further fusion. The fused results serve as the input of a ControlNet, which extracts multi-scale features for fine-grained injection of motion and identity. The framework is trained with weighted losses. We use a blend of real videos and image-simulated videos for training to compensate for the data scarcity.
@@ 3.1 Overview of Framework
The VideoAnydoor pipeline is illustrated in Fig. 2. To reconstruct the background within the masked region, we build our method on a 2D in-painting diffusion model. Following the latent diffusion model [31], we encode both the source video and the masked video with a VAE encoder to obtain the latent space representations $z_{ori}$ and $z_{mask}$. The corresponding masks are 8 times down-sampled as the mask latent. Subsequently, DDIM inversion [5] is applied to transform the clean latent $z_{ori}$ back to the noisy latent $z_{T}$. Then we concatenate $z_{T}$, $z_{mask}$ and $z_{mask}$ as a 9-channel tensor for 3D U-Net. To utilize the priors of video generation, we integrate the motion layers [11] into the in-painting model as the 3D U-Net to ensure essential temporal consistency. For coarse-grained control, we leverage the powerful visual encoder DINOv2 [25] as the ID extractor for ID preservation and use the bounding boxes as motion guidance. Before feeding the reference image into the extractor, we remove its background with a segmentor [16] and align the object to the image center to retain compact and ID-related representations. For fine-grained control, we adopt the interaction-friendly trajectory lines as the control signals and propose a pixel warper to warp the pixel details according to the desired motion for joint modeling of appearance details and precise motion. Finally, a weighted loss is used amplify the influence of key-points and design a novel image-video mix-training strategy to address the scarcity of high-quality video data.
For convenience, the trajectory map and correspondence reference image are denoted as $c_{mot}\in\mathcal{R}^{N\times 3\times H\times W}$ and $c_{ref-key}\in\mathcal{R}^{3\times H\times W}$ respectively. $V\in\mathcal{R}^{N\times 3\times H\times W}$, $A\in\mathcal{R}^{N\times 1\times H\times W}$ and $c_{ref}\in\mathcal{R}^{3\times H\times W}$ represent the original video, masks of the edited region and the reference image respectively. $N,H,W$ is the frame number, height, and width of the original video. The content and motion encoders are denoted as $E_{c}$ and $E_{m}$ respectively.
@@ 3.2 Pixel Warper
Weighted loss.
Inspired by  [39], we try to adopt a reweight scheme to improve the fine-grained modeling of identity and motion. Specifically, we solely perform reweighting on the region surrounding the trajectory. For trajectories with relatively large motion amplitudes, we apply greater weights to achieve more accurate motion control. Denote the ratio of the region that $i$-th trajectory covers as $R_{i}$, we perform 8 times down-sampling on the region, which is denoted as $\mathbf{A_{trj}^{i}}$. The weighted loss can be formulated as:
$\begin{split}\mathcal{L}=\sum_{i=1}^{N}(({\lambda R_{i}\mathbf{A_{trj}^{i}}+{(1-\mathbf{A_{trj}^{i}})/N}})\cdot\mathopen{}\mathclose{{}\left\|\delta-\delta^{*}}\right\|_{2}^{2}),\end{split}$ (2)
where $\lambda$ denotes the balancing loss weight. $\delta^{*},\delta$ are the prediction of the 3D U-Net and the target.
@@ 4.1 Experimental Setup
For comprehensive evaluation, we construct a benchmark consisting of around 200 videos collected from Pexel, which includes ten different categories (e.g., persons, dogs). We also make qualitative analysis on the ViViD [8] and CHDTF [54] test set to evaluate the performance for virtual video try-on and talking head generation.
@@ 4.3 Quantitative Comparison
Only Real-video Data 34.6 75.0 51.7
Only Static-image Data 33.9 73.8 51.2
FrozenDINOv2 33.5 74.3 51.6
w/o $\rm PixelWarper^{\dagger}$ 35.3 77.4 53.0
w/o Pixel Warper 33.8 72.4 48.5
w/o Weighted Loss 35.1 77.0 53.1
w/ Box Loss [39] 37.7 81.1 58.9
Ours-full 38.0 81.4 59.1
@@ 4.4 Ablation Studies
Only Real-video Data 66.5 67.5 69.9
Only Static-image Data 71.4 72.0 74.3
@@
w/o $\rm PixelWarper^{\dagger}$ 80.3 81.3 84.4
w/o $\rm PixelWarper$ 78.5 81.7 83.7
w/o Weighted Loss 75.4 84.2 85.1
w/ Box Loss [39] 88.0 91.1 92.3
Ours-full 88.3 91.5 92.5
@@
We conduct an investigation of the core components on ID preservation. From Tab. 4, we can observe that training with fixed DINOv2 induces much inferior performance. Training only on videos suffers from severe accuracy degradation across all the metrics. Moreover, our pixel warper can effectively help inject the appearance details according to the motion. The results also show that the weighted loss is superior to the box loss [39] and can bring a performance boost to the baseline by making the model focus on the foreground regions. The best performance can be achieved by training with all the modules.
@@
Random X-Pose points 80.6 82.4 83.0
Grid points 82.4 83.3 85.0
w/o NMS 82.2 83.0 84.3
Tight box 83.4 85.6 86.3
Ours-full 88.3 91.5 92.5
@@ 4.5 More Applications
Talking head generation. Besides video try-on, we further apply our method to talking head generation, which requires more precise control of tiny movements and preservation of face identities. Specifically, we conduct evaluations on the CHDTF dataset [54] and use 16 points with the largest movement in the face landmarks as the initial key-points. As demonstrated in Fig. 8, VideoAnydoor could give satisfactory performance for this task as well.
@@ 5 Conclusion
In this paper, we present VideoAnydoor for end-to-end video object insertion with precise motion control. Specifically, it can effectively characterize the reference target with an ID extractor when trained with a combination of videos and high-quality images. Moreover, it can achieve smooth motion consistency and effective preservation of appearance details through the proposed pixel warper. Our VideoAnydoor has promising performance on both video object insertion task and diverse precise video editing applications, e.g., (1) object insertion, (2) virtual video try-on, (3) video face swapping, and (4) multi-region editing. Extensive qualitative and quantitative experimental results demonstrate its superiority over previous methods on the alignment of both motion and identity with the given control signals. It provides a universal solution for general region-to-region mapping tasks as well.